Rumolt Motors has 6363 million shares outstanding with a price of $ 49 per share. In​ addition, Rumolt has issued bonds with a total current market value of $ 3,335 million. Suppose​ Rumolt's equity cost of capital is 12 % and its debt cost of capital is 8%.

a. What is​ Rumolt's pretax weighted average cost of​ capital?

b. If​ Rumolt's corporate tax rate is 21%​, what is its​ after-tax weighted average cost of​ capital?
Answer #1

a.

value of equity=63*49=3087 million

What is​ Rumolt's pretax weighted average cost of​capital

=(3087/(3087+3335))*12%+(3335/(3335+3087))*8%

=9.92%

b.

If​ Rumolt's corporate tax rate is 21%​, what is its​after-tax weighted average cost of​ capital

=(3087/(3087+3335))*12%+(3335/(3335+3087))*8%*(1-21%)

=9.05%

the above is answer..